### v2

# Next-order asymptotics for the volume of Schatten balls

Abstract.
The volume of the unit balls of self-adjoint finite-dimensional Schatten $p$-classes of $n\times n$-matrices, $1\leq p\leq\infty$, is only known exactly for $p=2$ and $p=\infty$. We give an asymptotic expansion of the logarithmic volume to order $o(n)$ for all $p>1$. The proof rests on asymptotics for the partition function of $\beta$-ensembles due to Leblé and Serfaty [Invent. Math. 210(3):645–757, 2017]. Independently, the case $p\geq 2$ was obtained by Dworaczek Guera, Memin and Pain [arXiv:2511.05386]. In the complex case the asymptotic expansion is continued to order $o(1)$ for all $p\geq 1$.
Key words and phrases:
entropy; partition function; Schatten norm; Ullman distribution
2020 Mathematics Subject Classification:
52A23; 47B10 (Primary) 60B20; 82B21 (Secondary)
1. Introduction
The Schatten or Schatten-von Neumann class consists of all compact operators between two Hilbert spaces whose sequence of singular values belongs to the Lebesgue sequence space $\ell_{p}$. It can be seen a non-commutative version of $\ell_{p}$ and is named after Schatten and von Neumann Schatten (1946, 1950); Schatten and von Neumann (1946). We refer to Pietsch (2007); Pisier and Xu (2003) for an overview of their role in functional analysis and Banach space geometry.
In this article, we study the geometry of finite-dimensional Schatten classes and in particular their unit balls. These have been investigated as test cases for conjectures in asymptotic geometric analysis Dadoun et al. (2023); Dworaczek Guera et al. (2025); Guédon and Paouris (2007); König et al. (1998); Radke and Vritsiou (2020), in the context of low-rank matrix recovery and information-based complexity Candes and Recht (2012); Chávez-Domínguez and Kutzarova (2015); Hinrichs et al. (2017, 2021); Prochno and Strzelecki (2022) or in quantum information theory Aubrun et al. (2011); Wilde (2013). Let $\mathbb{F}_{1}=\mathbb{R}$, $\mathbb{F}_{2}=\mathbb{C}$ and $\mathbb{F}_{4}=\mathbb{H}$ and denote the singular values of a matrix $A\in\mathbb{F}_{\beta}^{n\times n}$, where $\beta\in\{1,2,4\}$, by
$s_{1}(A)\geq\cdots\geq s_{n}(A)\geq 0.$
For $1\leq p\leq\infty$, the $p$-Schatten norm of $A$ is
$\|A\|_{p}=\begin{cases}\Big(\sum_{j=1}^{n}s_{j}(A)^{p}\Big)^{1/p}&\colon p<\infty\\ s_{1}(A)&\colon p=\infty.\end{cases}$
The associated unit ball and its intersection with the subspace of self-adjoint matrices $\mathcal{H}_{n}(\mathbb{F}_{\beta})=\{A\in\mathbb{F}_{\beta}^{n\times n}\colon A^{*}=A\}$ are denoted by
$B_{p,\beta}^{n}=\{A\in\mathbb{F}_{\beta}^{n\times n}\colon\|A\|_{p}\leq 1\}\quad\text{and}\quad\mathbb{B}_{p,\beta}^{n}=B_{p,\beta}^{n}\cap\mathcal{H}_{n}(\mathbb{F}_{\beta}).$
For self-adjoint matrices the Schatten $p$-norm $\|A\|_{p}$ becomes a $p$-norm of eigenvalues. In case of non-self adjoint matrices one can also study the rectangular case, see e.g. Juhos et al. (2024).
We shall identify $\mathbb{F}_{\beta}^{n\times n}$ with $\mathbb{R}^{\beta n^{2}}$ and $\mathcal{H}_{n}(\mathbb{F}_{\beta})$ with $\mathbb{R}^{d_{n}}$, where
$d_{n}:=\beta\frac{n(n-1)}{2}+n.$ (1)
Then $B_{p,\beta}^{n}$ and $\mathbb{B}_{p,\beta}^{n}$ are convex bodies (i.e. compact convex sets with nonempty interior) in $\mathbb{R}^{\beta n^{2}}$ and $\mathbb{R}^{d_{n}}$, respectively. In the following, we study their volumes $\mathrm{vol}(B_{p,\beta}^{n})$ and $\mathrm{vol}(\mathbb{B}_{p,\beta}^{n})$. Our aim is to give more precise asymptotics of the volume of $\mathbb{B}_{p,\beta}^{n}$ as $n\to\infty$.
The exact volume of $B_{p,\beta}^{n}$ and $\mathbb{B}_{p,\beta}^{n}$ is only known in the cases $p=2$ and $p=\infty$. For $p=2$, the Schatten $2$-balls are Euclidean unit balls of corresponding dimension and thus
$\mathrm{vol}(B_{2,\beta}^{n})=\frac{\pi^{\beta n^{2}/2}}{\Gamma(1+\frac{\beta n^{2}}{2})}\qquad\text{and}\qquad\mathrm{vol}(\mathbb{B}_{2,\beta}^{n})=\frac{\pi^{d_{n}/2}}{\Gamma(1+\frac{d_{n}}{2})},$
where $\Gamma(x)=\int_{0}^{\infty}t^{x-1}e^{-t}\mathrm{d}t$ denotes the Gamma function. For $p=\infty$ it follows from Saint Raymond’s work Saint Raymond (1984) and Selberg’s integral formula Anderson (1991); Selberg (1944) (see also Remark 8 below) that
$\mathrm{vol}(B_{\infty,\beta}^{n})=\frac{\prod_{j=0}^{n-1}\Gamma(1+j\frac{\beta}{2})}{\prod_{j=n}^{2n-1}\Gamma(1+j\frac{\beta}{2})}\pi^{\beta n^{2}/2},$ (2)
and in the self-adjoint case it holds that
$\mathrm{vol}(\mathbb{B}_{\infty,\beta}^{n})=2^{d_{n}}(2\pi)^{\beta n(n-1)/4}\prod_{j=0}^{n-1}\frac{\Gamma(1+j\frac{\beta}{2})^{2}}{\Gamma(2+(n+j-1)\frac{\beta}{2})}.$ (3)
For general $1\leq p\leq\infty$, exact volumes are unknown and instead their asymptotic behavior has been studied. Saint Raymond Saint Raymond (1984) derived $A(p)>0$ such that, for $\beta\in\{1,2\}$,
$\lim_{n\to\infty}n^{1/2+1/p}\mathrm{vol}(B_{p,\beta}^{n})^{1/\beta n^{2}}=\sqrt{\frac{2\pi}{\beta}}e^{3/4}A(p)$ (4)
and determined $A(2)=e^{-1/4}$ and $A(\infty)=\frac{1}{2}$. Guédon and Paouris Guédon and Paouris (2007) showed that $\mathrm{vol}(B_{p,4}^{n})^{1/4n^{2}}$ and $\mathrm{vol}(\mathbb{B}_{p,\beta}^{n})^{1/d_{n}},\beta\in\{1,2,4\}$ are of the same order $n^{-1/2-1/p}$. Kabluchko, Prochno and Thäle Kabluchko et al. (2020a) determined
$A(p)=\frac{1}{2}\bigg(\frac{p\sqrt{\pi}\Gamma(\frac{p}{2})}{\sqrt{e}\Gamma(\frac{p+1}{2})}\bigg)^{1/p}$ (5)
and moreover showed in Kabluchko et al. (2020b) that, for $\beta\in\{1,2,4\}$,
$\lim_{n\to\infty}n^{1/2+1/p}\mathrm{vol}(\mathbb{B}_{p,\beta}^{n})^{1/d_{n}}=\sqrt{\frac{4\pi}{\beta}}e^{3/4}A(p).$ (6)
We remark that one way to see that $A(p)$ is decreasing in $p$ is to observe that the same is true for the normalized volume $\mathrm{vol}(n^{1/p}\mathbb{B}_{p,\beta}^{n})$. Recently, Dadoun, Fradelizi, Guédon and Zitt (Dadoun et al., 2023, Remark 4.8) showed (4) for $\beta=4$. Thus, the asymptotics in (4) and (6) hold in all cases $\beta\in\{1,2,4\}$ and $1\leq p\leq\infty$.
In case of $p=\infty$, the quantity $A(\infty)=\frac{1}{2}$ can be expressed in terms of the minimal logarithmic energy of probability measures supported in [1,1][-1,1] attained by the arcsine distribution, see Saff and Totik (1997). We refer to Beltrán et al. (2024) and Brauchart (2024) for related asymptotics for discrete minimal logarithmic energy.
In case of $p<\infty$, the quantity $A(p)$ in (5) can be expressed as
$A(p)=\frac{1}{2}\Big(\frac{p\sqrt{\pi}\Gamma(\frac{p}{2})}{\sqrt{e}\Gamma(\frac{p+1}{2})}\Big)^{1/p}=\frac{1}{2}\Big(\frac{1}{\sqrt{e}\alpha_{p}}\Big)^{1/p},$
where $\alpha_{p}=\int_{\mathbb{R}}|x|^{p}\mathrm{d}\mu_{p}(x)$ is the $p$-th absolute moment of the Ullman distribution $\mu_{p}$ which has density
$f_{p}(x)=\frac{p}{\pi}\int_{|x|}^{1}\frac{t^{p-1}}{\sqrt{t^{2}-x^{2}}}\mathrm{d}t,\qquad x\in[-1,1].$ (7)
The Ullman distribution originated in the theory of orthogonal polynomials on $\mathbb{R}$ with weight $e^{-c|x|^{p}}$ and $A(p)$ is strongly related to the asymptotics of the sequence of leading coefficients, see Saff and Totik (1997) and (Van Assche, 1987, Sec. 4.1).
Our main contribution is the following asymptotic expansion for the logarithmic volume $\ln\mathrm{vol}(\mathbb{B}_{p,\beta}^{n})$. Recall that the differential entropy of a probability measure $\mu$ on $\mathbb{R}$ is
$\mathrm{Ent}(\mu)=-\int_{\mathbb{R}}\ln(f(x))f(x)\mathrm{d}x,$
whenever $\mu$ admits a Lebesgue density $f$ and $\mathrm{Ent}(\mu)=$-$\infty$ otherwise.
Theorem 1.
Let $\beta\in\{1,2,4\}$. If $p>1$, then, as $n\to\infty$,
$\ln\mathrm{vol}(\mathbb{B}_{p,\beta}^{n})=$ $-\frac{\beta}{2}\Big(\frac{1}{2}+\frac{1}{p}\Big)n^{2}\ln n$
$+\frac{\beta}{2}\Big(\frac{1}{2}\ln\frac{4\pi}{\beta}+\frac{3}{4}+\ln A(p)\Big)n^{2}$
$-\Big(1-\frac{\beta}{2}\Big)\Big(\frac{1}{2}+\frac{1}{p}\Big)n\ln n$
$+\Big(1-\frac{\beta}{2}\Big)\Big(\frac{1}{2}\ln\frac{4}{\beta\pi}+\frac{1}{2}+\frac{1}{2p}+\ln A(p)+\mathrm{Ent}(\mu_{p})\Big)n+o(n).$ (8)
Moreover, if $\beta=2$ and $p\geq 1$, then
$\ln\mathrm{vol}(\mathbb{B}_{p,2}^{n})=-\Big(\frac{1}{2}+\frac{1}{p}\Big)n^{2}\ln n+\Big(\frac{1}{2}\ln 2\pi+\frac{3}{4}+\ln A(p)\Big)n^{2}-\ln n+M_{p}+o(1),$ (9)
where
$M_{p}=\frac{5}{12}\ln p+\frac{1}{12}\ln 2.$
In case of $p=2$ it is known that $\mathrm{Ent}(\mu_{2})=\ln\pi-\frac{1}{2}$ where $\mu_{2}$ is the semi-circle law. Otherwise the entropy $\mathrm{Ent}(\mu_{p})$ of the Ullman distribution appears to be unknown. Independently, the expansion in (1) was recently obtained by Dworaczek Guera, Memin and Pain Dworaczek Guera et al. (2025) in case of $p\geq 2$.
In addition to Theorem 1 we obtain for $p=\infty$ from the exact formula (3) and asymptotics for the Barnes $G$-function satisfying $G(n+1)=\prod_{k=0}^{n-1}\Gamma(1+k)$, $n\in\mathbb{N}$, that
$\ln\mathrm{vol}(\mathbb{B}_{\infty,2}^{n})=$ $-\frac{1}{2}n^{2}\ln n+\Big(\frac{1}{2}\ln 2\pi+\frac{3}{4}+\ln A(\infty)\Big)n^{2}$
$-\frac{1}{6}\ln n+\frac{1}{12}\ln 2+2\zeta^{\prime}(-1)+o(1),$ (10)
where $\zeta$ is the Riemann zeta function and $\zeta^{\prime}(-1)=\frac{1}{12}-\ln A$ with $A=1.282$... being the Glaisher-Kinkelin constant. The logarithmic term is what we would expect from
$\ln\mathrm{vol}(\mathbb{B}_{p,\beta}^{n})\leq\ln\mathrm{vol}(\mathbb{B}_{\infty,\beta}^{n})\leq\ln\mathrm{vol}(\mathbb{B}_{p,\beta}^{n})+\frac{d_{n}}{p}\ln n,$
upon setting $p=n^{2}(\ln n)^{2}$ such that the difference between the logarithmic volumes vanishes. The discrepancy in the constant terms in (9) and (1) is because the $o(1)$-term in (9) implicitly depends on $p$.
The proof of Theorem 1 relies on the fact that the volume of $\mathbb{B}_{p,\beta}^{n}$ can be expressed in terms of the integral
$Z_{n,p,\beta}=\int_{\mathbb{R}^{n}}\prod_{1\leq i<j\leq n}|x_{i}-x_{j}|^{\beta}\prod_{i=1}^{n}e^{-\frac{\beta}{2}n\,v_{p}|x_{i}|^{p}}\mathrm{d}x,$ (11)
where
$v_{p}=\frac{\sqrt{\pi}\Gamma(\frac{p}{2})}{\Gamma(\frac{p+1}{2})}=\frac{1}{p\alpha_{p}}$
is chosen for normalization purposes. The quantity $Z_{n,p,\beta}$ is the partition function of a one-dimensional $\beta$-ensemble with potential $V(x)=v_{p}|x|^{p}$, where eigenvalues/particles are distributed with joint density
$f_{n,p,\beta}(x)=\frac{1}{Z_{n,p,\beta}}\prod_{1\leq i<j\leq n}|x_{i}-x_{j}|^{\beta}\prod_{i=1}^{n}e^{-\frac{\beta}{2}n\,v_{p}|x_{i}|^{p}},\qquad x\in\mathbb{R}^{n}.$ (12)
The parameter $\beta$ stands for the inverse temperature and may take any positive value, while for the matrix representation we restrict to $\beta\in\{1,2,4\}$. In case of $p=2$ the integral $Z_{n,2,\beta}$ is known exactly via the Mehta-Selberg integral. We refer to Anderson et al. (2010) and Forrester (2010) for more information on $\beta$-ensembles.
In this context the Ullman distribution arises as the corresponding equilibrium measure minimizing the functional
$\mu\mapsto I_{p}(\mu)=-\int_{\mathbb{R}}\int_{\mathbb{R}}\ln|x-y|\mathrm{d}\mu(x)\mathrm{d}\mu(y)+v_{p}\int_{\mathbb{R}}|x|^{p}\mathrm{d}\mu(x)$
over $\mu\in\mathcal{P}(\mathbb{R})$, where $\mathcal{P}(\mathbb{R})$ denotes the space of probability measures on $\mathbb{R}$. More precisely, if $X=(X_{1},\dots,X_{n})$ is distributed according to the density $f_{n,p,\beta}$, then the random empirical measure $\mu_{n,p}=\frac{1}{n}\sum_{i=1}^{n}\delta_{X_{i}}$ tends to $\mu_{p}$ as $n\to\infty$. Moreover, large deviation principles due to Ben Arous and Guionnet (1997) and Hiai and Petz (2000) yield
$\lim_{n\to\infty}\frac{2}{\beta n^{2}}\ln Z_{n,p,\beta}=I_{p}(\mu_{p})=\ln 2+\frac{3}{2p}.$ (13)
Earlier work De Monvel et al. (1995); Johansson (1998) also identified the limiting constant but with different techniques. In Dadoun et al. (2023), the limit (13) is used to prove (6). The large deviation principle for the corresponding tagged empirical fields due to Leblé and Serfaty Leblé and Serfaty (2017) allows to refine (13) as follows.
Theorem 2.
Let $\beta>0$ and $p>1$. As $n\to\infty$,
$\ln Z_{n,p,\beta}=-\frac{\beta}{2}n^{2}I_{p}(\mu_{p})+\frac{\beta}{2}n\ln n-C(\beta)n+\Big(1-\frac{\beta}{2}\Big)\mathrm{Ent}(\mu_{p})n+o(n),$
where
$C(\beta)=\Big(1-\frac{\beta}{2}\Big)\ln\frac{\beta}{2}-\frac{\beta}{2}\ln 2\pi+\frac{\beta}{2}+\ln\Gamma\Big(\frac{\beta}{2}\Big).$ (14)
Theorem 2 forms the core of the proof of (1) in Theorem 1. Independently, in the recent work Dworaczek Guera et al. (2025) the authors deduced the statement of Theorem 2 in case of $p\geq 2$ from a central limit theorem for linear eigenvalue statistics. The restriction to $p\geq 2$ results from regularity requirements of the underlying approach, see also Dadoun et al. (2023). The milder restriction to $p>1$ in Theorem 2 is due to the fact that in this case the Ullman density $f_{p}$ is Hölder continuous inside its support which is already sufficient for the approach in Leblé and Serfaty (2017).
The term $\mathrm{Ent}(\mu_{p})$ in Theorem 2 arises via a rescaling property of the specific entropy which is one part of the rate function in the large deviation principle in Leblé and Serfaty (2017). In case of $\beta=2$, the contribution of $\mathrm{Ent}(\mu_{p})$ vanishes and more precise asymptotics hold. The following is a consequence of (Claeys et al., 2023, Prop. 1.1), derived from Kriecherbauer and McLaughlin (1999), and (Claeys et al., 2023, Thm. 1.9).
Theorem 3.
Let $p\geq 1$. As $n\to\infty$,
$\ln Z_{n,p,2}=-n^{2}I_{p}(\mu_{p})+n\ln n+(\ln 2\pi-1)n+\frac{5}{12}\ln n-\frac{1}{12}\ln\frac{p}{2}+\zeta^{\prime}(-1)+o(1),$
where $\zeta$ is the Riemann zeta function.
Theorem 3 combined with Lemma 7 and Lemma 9 below yields (9) in Theorem 1. In case of $p$ an even integer, the potential $V(x)=v_{p}|x|^{p}$ is an analytic function and (Borot and Guionnet, 2013, Prop. 1.2) yields a full asymptotic expansion of $\ln Z_{n,p,\beta}$, and consequently of $\ln\mathrm{vol}(\mathbb{B}_{p,\beta}^{n})$, with coefficients defined in terms of integrals of the Stieltjes transform of $\mu_{p}$, see also (Borot and Guionnet, 2024, Sec. 7) and Shcherbina (2013, 2014).
In the case of non self-adjoint matrices, for $\beta\in\{1,2,4\}$ and $1\leq p<\infty$, the volume of the unit ball $B_{p,\beta}^{n}$ may be written in terms of
$\tilde{Z}_{n,p,\beta}=\int_{(\mathbb{R}_{+})^{n}}\prod_{1\leq i<j\leq n}|x_{i}-x_{j}|^{\beta}\prod_{i=1}^{n}x_{i}^{\beta/2-1}\prod_{i=1}^{n}e^{-\beta n\,v_{p}x_{i}^{p/2}}\mathrm{d}x,$
see e.g. (Dadoun et al., 2023, Sec. 4.2). This corresponds to the partition function of a Laguerre $\beta$-ensemble with potential $V(x)=v_{p}|x|^{p/2}$. The leading order asymptotics for $\tilde{Z}_{n,p,\beta}$ can again be obtained via large deviations, leading to (4). However, the results in Leblé and Serfaty (2017) do not apply because of a blow-up of the corresponding equilibrium density near the boundary of its support. It goes beyond the scope of the present work to derive next-order asymptotics for $\tilde{Z}_{n,p,\beta}$ and $B_{p,\beta}^{n}$. For completeness, we note that in case of $p=\infty$ it can be derived that (see e.g. Juhos et al. (2024))
$\ln\mathrm{vol}(B_{\infty,\beta}^{n})=$ $-\frac{\beta}{2}n^{2}\ln n+\beta\Big(\frac{1}{2}\ln\frac{2\pi}{\beta}+\frac{3}{4}+\ln A(\infty)\Big)n^{2}$
$+\Big(1-\frac{\beta}{2}\Big)\ln(A(\infty))n+\frac{-3+\frac{\beta}{2}+\frac{2}{\beta}}{12}\ln n+O(1).$
The unit balls $B_{p,\beta}^{n}$ and $\mathbb{B}_{p,\beta}^{n}$ can even be studied for $0<p<1$. Then (4) and (6) as well as Theorem 3 (with a possibly modified constant term) continue to hold. However, the proof of Theorem 1 and the underlying large deviation principle in Leblé and Serfaty (2017) appear to be limited to $p\geq 1$. This is because the rate function in Leblé and Serfaty (2017) is the free energy which is uniquely minimized by the sineβ-process, see Erbar et al. (2021); Huesmann and Müller (2025). While the microscopic limit point process in the bulk of many $\beta$-ensembles is given by the universal sineβ-process, see Bourgade et al. (2014); Mehta (1991); Valkó and Virág (2009), for $\beta=2$ and $p<1$ the density $f_{p}$ develops a singularity at zero near which the limit point process depends on $p$, see Claeys et al. (2023). In case of $\beta\in\{1,4\}$, it is an open problem to extend Theorem 1 to $p=1$ for which
$f_{1}(x)=\frac{1}{\pi}\ln\frac{1+\sqrt{1-x^{2}}}{|x|},\qquad x\in[-1,1].$
The logarithmic singularity at the origin prevents direct application of the results in Leblé and Serfaty (2017) but possibly behaves well enough such that the approach may be adapted to this case.
Remark 4.
Heuristically, the unit balls $B_{p,\beta}^{n}$ and $\mathbb{B}_{p,\beta}^{n}$ are non-commutative versions of $\mathbb{B}_{p}^{n}=\{x\in\mathbb{R}^{n}\colon\sum_{i=1}^{n}|x_{i}|^{p}\leq 1\}$. To compare, note that
$\mathrm{vol}(n^{1/p}\mathbb{B}_{p}^{n})=n^{n/p}\frac{2^{n}\Gamma(1+\frac{1}{p})^{n}}{\Gamma(1+\frac{n}{p})}=\frac{1+o(1)}{\sqrt{2\pi n/p}}e^{\mathrm{Ent}(\nu_{p})n},$
where $\nu_{p}$ has density proportional to $e^{-|x|^{p}/p}$ and maximizes entropy over $\mu\in\mathcal{P}(\mathbb{R})$ with $\int|x|^{p}\mathrm{d}\mu(x)\leq 1$, see also Kabluchko and Prochno (2021). Similarly, the Ullman distribution $\mu_{p}$ maximizes free entropy over $\mu\in\mathcal{P}(\mathbb{R})$ with $\int|x|^{p}\mathrm{d}\mu(x)\leq\alpha_{p}$, see (Hiai and Petz, 2000, Prop. 5.34). In case of $p=\infty$, one can interpret $\nu_{\infty}$ as the uniform distribution on [1,1][-1,1], which maximizes entropy, and $\mu_{\infty}$ as the arcsine distribution, which maximizes free entropy. Formally, see Lemma 7 below, the case $\beta=0$ corresponds to the commutative case where a Poisson point process appears in the microscopic limit, see also Allez and Dumaz (2014); Benaych-Georges and Péché (2015); Leblé (2016). Possibly, a careful analysis of the behavior of the asymptotics for $\ln Z_{n,p,\beta}$ as $\beta\to 0$ could allow a computation of $\mathrm{Ent}(\mu_{p})$ in terms of $\mathrm{Ent}(\nu_{p})$.
Remark 5.
Theorem 1 can be combined with the stochastic representation in (Kabluchko et al., 2020b, Cor. 4.3) of a random matrix drawn uniformly from a Schatten ball $\mathbb{B}_{p,\beta}^{n}$ to obtain information on critical intersections of Schatten $p$-balls in the spirit of Kabluchko et al. (2020b) and Sonnleitner and Thäle (2025).
2. Proofs
In the following, we provide the proofs of Theorems 1 and 2. First, in Lemma 6 we prove Hölder regularity of the Ullman density. Then Theorem 2 is a direct consequence of Leblé and Serfaty (2017). Subsequently, in Lemma 7 we recall a Weyl type integration formula providing the link between $\mathrm{vol}(\mathbb{B}_{p,\beta}^{n})$ and $Z_{n,p,\beta}$. Finally, we deduce Theorem 1 from Theorem 2 with the help of a known asymptotic expansion in Lemma 9.
Corollary 1.1 in Leblé and Serfaty (2017) states that the asymptotics in Theorem 2 hold, i.e.
$\ln Z_{n,p,\beta}=-\frac{\beta}{2}n^{2}I_{p}(\mu_{p})+\frac{\beta}{2}n\ln n-C(\beta)n+\Big(1-\frac{\beta}{2}\Big)\mathrm{Ent}(\mu_{p})n+o(n),$ (15)
provided that the function $V(x)=v_{p}|x|^{p}$ satisfies certain assumptions (H1)-(H5). These involve regularity of the density of the equilibrium measure which in this case is the Ullman density
$f_{p}(x)=\frac{p}{\pi}\int_{|x|}^{1}\frac{t^{p-1}}{\sqrt{t^{2}-x^{2}}}\mathrm{d}t,\qquad x\in[-1,1].$
It is known from (Deift et al., 1998, Thm. 1.34) that if $V\in C^{3}$, that is if $p\geq 3$, then the density of the equilibrium measure is Hölder continuous of order $\frac{1}{2}$ and behaves like $\sqrt{1-|x|}$ near the boundary of its support. To include smaller values of $p$, we verify the following statement directly.
Lemma 6.
Let $p>1$. Then $f_{p}$ is Hölder continuous of order $\min\{p-1,\frac{1}{2}\}$ on [1,1][-1,1]. Moreover, it is Hölder continuous of order $\frac{1}{2}$ in neighborhoods of 1-1 and $1$ and there exist positive constants $c$ and $C$ such that
$c\sqrt{1-|x|}\leq f_{p}(x)\leq C\sqrt{1-|x|}\qquad\text{for}\qquad|x|\geq\frac{1}{2}.$ (16)
Proof.
First, we verify the behavior of $f_{p}$ near the boundary points. By a change of variables, for $x\in(-1,1)\setminus\{0\}$,
$\frac{\pi}{p}f_{p}(x)=\int_{|x|}^{1}\frac{t^{p-1}}{\sqrt{t^{2}-x^{2}}}\mathrm{d}t=|x|^{p-1}\int_{0}^{\arccos|x|}\frac{1}{\cos^{p}\alpha}\mathrm{d}\alpha.$ (17)
Combining (17) with the asymptotics $\arccos|x|=\sqrt{2(1-|x|)}+o(1-|x|)$ as $|x|\to 1$ proves (16) which in turn implies that $f_{p}$ is Hölder continuous of order $\frac{1}{2}$ at the boundary points $\pm 1$.
Next, we show that $f_{p}$ is in fact continuously differentiable at $x\in(-1,1)\setminus\{0\}$. For symmetry reasons it is enough to consider $x\in(0,1)$. Let $0<|h|<\min\{x,1-x\}$ and write
$\frac{\pi}{p}\big(f_{p}(x+h)-f_{p}(x)\big)=$ $\big((x+h)^{p-1}-x^{p-1}\big)\int_{0}^{\arccos x}\frac{1}{\cos^{p}\alpha}\mathrm{d}\alpha$
$-(x+h)^{p-1}\int_{\arccos(x+h)}^{\arccos x}\frac{1}{\cos^{p}\alpha}\mathrm{d}\alpha.$
Using $\arccos(x+h)=\arccos x-\frac{h}{\sqrt{1-x^{2}}}+o(h)$ and letting $h\to 0$ shows that $f_{p}$ is continuously differentiable at $x$ with derivative
$f_{p}^{\prime}(x)=\frac{p-1}{x}f_{p}(x)-\frac{p}{\pi}\frac{1}{x\sqrt{1-x^{2}}}.$
It remains to check Hölder continuity at zero. Let $h\in(0,1)$ and write
$\frac{\pi}{p}\big(f_{p}(h)-f_{p}(0)\big)$ $=\int_{h}^{1}\frac{t^{p-1}}{\sqrt{t^{2}-h^{2}}}\mathrm{d}t-\int_{0}^{1}t^{p-2}\mathrm{d}t$
$=\int_{h}^{1}t^{p-1}\Big(\frac{1}{\sqrt{t^{2}-h^{2}}}-\frac{1}{t}\Big)\mathrm{d}t-\frac{h^{p-1}}{p-1}.$
The latter integral can be estimated by
$\int_{h}^{1}t^{p-1}\Big(\frac{1}{\sqrt{t^{2}-h^{2}}}-\frac{1}{t}\Big)\mathrm{d}t$ $\leq h^{2}\int_{h}^{1}\frac{t^{p-3}}{\sqrt{t^{2}-h^{2}}}\mathrm{d}t$
$=h^{p-1}\int_{1}^{1/h}\frac{u^{p-3}}{\sqrt{u^{2}-1}}\mathrm{d}u.$
In particular, if $1<p<2$, then
$0\leq\frac{\pi}{p}\frac{f_{p}(h)-f_{p}(0)}{h^{p-1}}\leq\int_{1}^{\infty}\frac{u^{p-3}}{\sqrt{u^{2}-1}}\mathrm{d}u<\infty.$ (18)
Thus, in this case we obtain that $f_{p}$ is Hölder continuous of order p1p-1 at 0. For $p\geq 2$, a similar argument shows that $f_{p}$ is continuously differentiable at 0 with $f_{p}^{\prime}(0)=0$. In conclusion, if $p>1$, the function $f_{p}$ is Hölder continuous of order $\min\{p-1,\frac{1}{2}\}$ on [1,1][-1,1]. This completes the proof. ∎
We can now deduce Theorem 2.
Proof of Theorem 2.
Let $p>1$. In order to apply (Leblé and Serfaty, 2017, Cor. 1.1) to $V(x)=v_{p}|x|^{p}$, we verify assumptions (H1)-(H5) there. The function $V$ is lower semi-continuous, bounded from below, everywhere finite and grows polynomially as $|x|\to\infty$. Thus, assumptions (H1) and (H2) in Leblé and Serfaty (2017) are satisfied. By Lemma 6 the density $f_{p}$ of the equilibrium measure $\mu_{p}$ is Hölder continuous of order $\min\{p-1,\frac{1}{2}\}$ on its support $\Sigma=[-1,1]$ (and Hölder continuous of order $\frac{1}{2}$ away from zero). Thus, assumption (H3) is satisfied. For assumption (H4) we note that $\Sigma$ is connected and its boundary is given by $\{-1,1\}$. Further, by Lemma 6, the density $f_{p}$ behaves like $c\sqrt{1-|x|}$ in neighborhoods of the boundary points and is Hölder continuous of order $\frac{1}{2}$ there. Thus, assumption (H4) is satisfied. Assumption (H5) is satisfied since $x\mapsto\exp(-c(|x|^{p}-\ln|x|))$ is integrable on $\mathbb{R}$ for every $c>0$. Thus, we can apply (Leblé and Serfaty, 2017, Cor. 1.1) to obtain (15). Finally, the constant $C(\beta)$ can be obtained from the known full expansion of the partition function in the Gaussian case $p=2$, see e.g. (Borot and Guionnet, 2024, Sec. 7). ∎
For the proof of Theorem 1 we provide the following well-known link between Schatten $p$-balls and $\beta$-ensembles. It is a combination of a Weyl type integration formula, see Anderson et al. (2010); Saint Raymond (1984), with a classical trick in convexity going back to Dirichlet (1839) for computing the volume of the unit ball of $\ell_{p}^{n}$, see Dadoun et al. (2023); Guédon and Paouris (2007); König et al. (1998). Recall that $d_{n}=\frac{\beta}{2}n^{2}+(1-\frac{\beta}{2})n$ and $\mathbb{B}_{p}^{n}=\{x\in\mathbb{R}^{n}\colon\sum_{i=1}^{n}|x_{i}|^{p}\leq 1\}$.
Lemma 7.
Let $\beta\in\{1,2,4\}$ and $1\leq p<\infty$. Then
$\mathrm{vol}(\mathbb{B}_{p,\beta}^{n})$ $=c_{n}\int_{\mathbb{B}_{p}^{n}}\prod_{1\leq i<j\leq n}|x_{i}-x_{j}|^{\beta}\mathrm{d}x$
$=\frac{c_{n}}{\Gamma(1+\frac{d_{n}}{p})}\Big(\frac{n\beta v_{p}}{2}\Big)^{d_{n}/p}Z_{n,p,\beta}$
where $Z_{n,p,\beta}$ and $v_{p}$ are as in (11) and
$c_{n}=\frac{1}{n!}\bigg(\frac{\Gamma(\frac{\beta}{2})}{(2\pi)^{\beta/2}}\bigg)^{n}\prod_{k=1}^{n}\frac{(2\pi)^{\beta k/2}}{\Gamma(\frac{\beta k}{2})}.$
Remark 8.
For $p=\infty$, the statement of Lemma 7 becomes
$\mathrm{vol}(\mathbb{B}_{\infty,\beta}^{n})=c_{n}\int_{[-1,1]^{n}}\prod_{1\leq i<j\leq n}|x_{i}-x_{j}|^{\beta}\mathrm{d}x.$
Then (3) is a consequence of Selberg’s integral formula.
As a final ingredient, we shall need the following well-known asymptotics for $c_{n}$ as in Lemma 7.
Lemma 9.
Let $\beta\in\{1,2,4\}$. Then, as $n\to\infty$,
$\ln c_{n}=$ $-\frac{\beta}{4}n^{2}\ln n+\frac{\beta}{2}\Big(\frac{1}{2}\ln\frac{4\pi}{\beta}+\frac{3}{4}\Big)n^{2}-\frac{1}{2}\Big(1+\frac{\beta}{2}\Big)n\ln n$
$+\Big(\frac{1}{2}\Big(1-\frac{\beta}{2}\Big)\ln\beta-\frac{1}{2}\Big(1+\frac{\beta}{2}\Big)(\ln\pi-1)-\ln 2+\ln\Gamma\Big(\frac{\beta}{2}\Big)\Big)n$
$-\frac{3+\frac{\beta}{2}+\frac{2}{\beta}}{12}\ln n+a_{\beta}+o(1),$
where $a_{\beta}\in\mathbb{R}$ is some constant with $a_{2}=\frac{1}{2}\ln 2\pi-\zeta^{\prime}(-1)$ with $\zeta$ being the Riemann zeta function.
Proof.
It follows from the definition of $c_{n}$ that
$c_{n}=\frac{1}{n!}\Gamma_{2}\Big(n+1;\frac{2}{\beta},1\Big)\bigg(\frac{\Gamma(\frac{\beta}{2})}{(\frac{\beta}{2})^{\frac{\beta}{2}}}\bigg)^{n}\Big(\frac{4\pi}{\beta}\Big)^{\frac{\beta}{4}n^{2}-(\frac{1}{2}+\frac{\beta}{4})n},$
where $\Gamma_{2}(x;\frac{2}{\beta},1)$ is a particular instance of the Barnes double Gamma function, see Borot and Guionnet (2024); Spreafico (2009). Adapting (Borot and Guionnet, 2024, eq. (7.17)), the following asymptotic expansion holds
$\ln\Gamma_{2}\Big(n+1;\frac{2}{\beta},1\Big)=$ $-\frac{\beta}{4}n^{2}\ln n+\frac{3\beta}{8}n^{2}+\frac{1}{2}\Big(1-\frac{\beta}{2}\Big)n\ln n-\frac{1}{2}\Big(1-\frac{\beta}{2}\Big)n$
$+\bigg(\frac{1}{2}-\frac{3+\frac{\beta}{2}+\frac{2}{\beta}}{12}\bigg)\ln n+a_{\beta}^{\prime}+o(1),$
where $a_{\beta}^{\prime}\in\mathbb{R}$ is some constant with $a_{2}^{\prime}=\frac{1}{2}\ln 2\pi-\zeta^{\prime}(-1)$. Combined with Stirling’s approximation $\ln n!=n\ln n-n+\frac{1}{2}\ln(2\pi n)+o(1)$ we deduce the claimed expansion. ∎
We are now ready to deduce the asymptotics for the volume of the self-adjoint Schatten classes $\mathbb{B}_{p,\beta}^{n}$ given in Theorem 1.
Proof of Theorem 1.
Let $\beta\in\{1,2,4\}$ and $p>1$. Recall that $d_{n}=\frac{\beta}{2}n^{2}+(1-\frac{\beta}{2})n$. By Lemma 7 it holds that
$\ln\mathrm{vol}(\mathbb{B}_{p,\beta}^{n})=\ln c_{n}+\frac{d_{n}}{p}\ln\Big(\frac{\beta nv_{p}}{2}\Big)-\ln\Gamma\Big(1+\frac{d_{n}}{p}\Big)+\ln Z_{n,p,\beta}.$ (19)
By Stirling’s approximation, as $n\to\infty$,
$\ln\Gamma\Big(1+\frac{d_{n}}{p}\Big)$ $=\frac{d_{n}}{p}\Big(\ln\frac{d_{n}}{p}-1\Big)+\frac{1}{2}\ln\Big(\frac{2\pi d_{n}}{p}\Big)+o(1)$
$=\frac{d_{n}}{p}\Big(2\ln n+\ln\frac{\beta}{2p}-1+\Big(1-\frac{\beta}{2}\Big)\frac{2}{\beta n}\Big)+\frac{1}{2}\ln\Big(\frac{\pi\beta n^{2}}{p}\Big)+o(1).$ (20)
Therefore,
$\frac{d_{n}}{p}\ln\Big(\frac{\beta nv_{p}}{2}\Big)-\ln\Gamma\Big(1+\frac{d_{n}}{p}\Big)=$ $-\frac{d_{n}}{p}\Big(\ln n-\ln(pv_{p})-1+\Big(1-\frac{\beta}{2}\Big)\frac{2}{\beta n}\Big)+o(n)$
$=$ $-\frac{\beta}{2p}n^{2}\ln n+\frac{\beta}{2p}\big(\ln(pv_{p})+1\big)n^{2}$
$-\frac{1}{p}\Big(1-\frac{\beta}{2}\Big)n\ln n+\frac{\ln(pv_{p})}{p}\Big(1-\frac{\beta}{2}\Big)n+o(n).$ (21)
Combining Theorem 2, Lemma 9 and (2) with (19) and collecting terms completes the proof of (1) in Theorem 1.
For the proof of (9) in Theorem 1 we proceed analogously, replacing Theorem 2 by Theorem 3. ∎
Acknowledgement
The author is grateful to Thomas Leblé for a helpful discussion regarding the case $p=1$. Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the SPP 2265 Random Geometric Systems and under Germany’s Excellence Strategy EXC 2044/2 - 390685587, Mathematics Münster: Dynamics-Geometry-Structure. This research was funded in whole or in part by the Austrian Science Fund (FWF) [Grant DOI: 10.55776/J4777]. For open access purposes, the author has applied a CC BY public copyright license to any author-accepted manuscript version arising from this submission.